# Application of AI-based Models for Online Fraud Detection and Analysis

### v1

@@ Abstract
Background Fraud is a prevalent offence that extends beyond financial loss, impacting victims emotionally, psychologically, and physically. The advancements in online communication technologies, allow for online fraud to thrive in this vast network, with fraudsters increasingly using these channels for deception. With the progression of technologies like Generative Artificial Intelligence (GenAI), there is a growing concern that fraud will increase and scale, using these advanced methods such as deep-fakes in phishing campaigns. However, the application of AI in detecting and analyzing online fraud remains understudied. This review addresses this gap by investigating AI’s role in analyzing online fraud using text data.
@@
Methods We conducted a Systematic Literature Review (SLR) on AI and Natural Language Processing (NLP) techniques for online fraud detection. The review adhered to the PRISMA-ScR protocol, with eligibility criteria including language, publication type, relevance to online fraud, use of text data, and AI methodologies. Out of $2,457$ academic records screened, $350$ met our eligibility criteria, and $223$ were analyzed and included herein.
@@
Results We report the state-of-the-art NLP techniques used to analyse various online fraud categories; the data sources used for training the NLP models; the NLP algorithms and models built; and the performance metrics employed for model evaluation. We find that the current state of research on online fraud is broken into the various scam activities that take place, and more specifically, we identify $16$ different frauds that researchers focus on. Finally, we present the latest and best performing AI methods used towards detecting online scam and fraud activities.
@@
Conclusions This SLR enhances the academic understanding of AI-based detection methods for online fraud and offers insights for policymakers, law enforcement, and businesses on safeguarding against such activities. We conclude that focusing on specific scams lacks generalization, as multiple models are required for different fraud types. Furthermore, the evolving nature of scams limits the effectiveness of models trained on outdated data. We also identify issues in data limitation or training bias reporting from researchers, along with issues on model performance reporting, with some studies selectively presenting metrics, leading to potential biases in model evaluation.
@@
*NOTE: This manuscript is currently under peer review. Please be aware that the content may be subject to revision in future versions.
@@
Online fraud has emerged as one of the most pervasive and challenging threats in the digital age affecting individuals of all ages, businesses of different sizes, and governments. Defined broadly, online fraud is an umbrella term that involves acts of deception or deliberate impersonation, taking place on the Internet, towards the personal gain of the fraudster, often resulting in a financial loss for the victim ukfinance2023 . Other than financial losses, fraud can have a wide range of impacts on victims. These include emotional and psychological effects such as anger, fear, shame, depression, decrease in confidence, and trauma; impacts on physical and mental well-being; harms relationships and leads to loneliness and isolation; and causes negative changes in behaviour parliament_post_2020 . Although there is evidence suggesting that certain socio-demographic groups face higher risks of fraud–such as women aged 25-34 or 35-44 experiencing fraud at rates of $7.6\%$, and individuals in the highest income bracket at $8.3\%$–fraud is impacting individuals across all demographics parliament_post_2020 and sometimes in different ways. For example, victims earning £20,000 or less, those aged 65 and over, and female victims, reported impacts on their self-confidence, at rates of $46\%$, $43\%$, and $41\%$, respectively, compared to the overall victim rate of $35\%$.
@@
For the year ending March 2023, the Crime Survey for England and Wales estimated 3.5 million fraud offences including online fraud  ons2023 . For example, advance fee fraud increased significantly highly from 60,000 to 391,000 offences compared to the year ending March 2020. This increase is largely due to society’s growing reliance on the Internet and digital platforms for everyday services, transactions and communications. According to The Office of Communication (Ofcom) ofcom2023 , $92\%$ of adults in the UK use the Internet for a wide variety of activities including communication, education, and entertainment. Activities such as banking, shopping, and socialising are increasingly happening via online platforms, expanding the landscape for fraudsters to exploit vulnerabilities or use these platforms to deceive victims. In 2020, online shopping scams made up $38\%$ of all reported scams worldwide, an increase of $6\%$ compared to pre-Covid-19 outbreak statista-ecommerce-fraud .
@@
Online fraud encompasses a wide range of deceptive activities, including identity theft, phishing, advance fee fraud, romance scams, fraudulent investment scams, and more. It is important to highlight that there is no universally accepted definition of “online fraud,” and the term is often used interchangeably with the term “scam.” Legally, “fraud is defined as false representation to cause loss to another or to expose another to a risk of loss” ukfraudlegislation2006 , and scam is the process where criminals gain the trust of victims to deceive or cheat victims ukscamglossary2006 , through false representation and other means, so that the victim trust them, which in turn results in various kinds of losses for the victim. The National Fraud Authority published a literature review fraudtypologiesuk where they compared the distinction of the term fraud as defined by the amended Fraud Act 2006 ukfraudlegislation2006 and the typology produced by Levi levi2008definingfraud . They found fraud embraces a broad scope of different crimes, whereas scams often are focused on fraud against individuals and small firms. For example, different scams like advance fee, romance, tech support, etc., all fall under the fraud umbrella, but they are also deception methods, which are in part scams. Hence, in this SLR, we use both terms as the various scams are the different deception methods scammers use to trick victims, and fraud is the term that includes all scams.
@@
Online frauds exploit the virtual nature of the Internet and the anonymity it provides to reach out to victims. This virtual environment, coupled with jurisdictional challenges (where offenders and victims may be in different regions of the world), makes fraud difficult to detect and prevent using traditional policing techniques. The complexity of online fraud is further heightened by its evolving nature, as fraudsters continuously adapt their techniques to bypass new security measures and exploit emerging technologies to target on new victims skidmore24 .
@@
Given the scale and impact of online fraud, there is a need for new methods to detect and prevent such activities. The use of NLP, in combination with other AI, has been proposed for identifying, characterising, and detecting fraudulent patterns in applications like phishing 9795286 , fake job advertisement amaar2022detection and analysing scam patterns lwin2023supporting which could help develop preventive measures and mitigate risks of online fraud. However, understanding the current state of AI techniques in combating online fraud, the data sources used, the evaluation methods for AI models, and the specific types of fraud that are most prevalent remain a significant challenge. This is due to the constant emergence of new fraud activities that use various communication mediums and social engineering attacks, in an attempt by fraudsters to remain undetected. Therefore, there is a pressing need to shift from detecting and analysing the effects of fraud to early detecting emerging fraudulent activities online and new methods of social engineering.
@@
This systematic literature review (SLR) aims to address these challenges by providing a comprehensive overview of the state-of-the-art AI techniques used to detect fraudulent activities online. More specifically, we explore the data sources widely used by researchers to study online fraud, how researchers evaluate the methodologies employed to assess AI models, and identify the most popular types of online fraud targeted by researchers. By synthesising findings from academic papers, this review aims to provide a thorough understanding of the current landscape of online fraud detection and prevention, highlighting gaps in existing research, and proposing directions for future studies.
Research Questions. The research question for the Systematic Literaure Review (SLR) is as follows:
  • RQ2: What are the data sources that researchers use to analyse online fraud?
  • RQ3: How do researchers evaluate their AI models?
  • RQ4: What are the most popular fraud activities that researchers study that fraudsters use to approach their potential victims?
Although a wide number of studies have explored the application of AI for fraud detection and other types of cybercrime, we are not aware of any systematic literature reviews that have examined the application of AI models using text data. The focus of this SLR is AI-based models that study textual data to detect and gain insights related to online fraud. Thus, the focus of this study is to identify Natural Language Processing models that are used to detect online fraud.
@@ 1 Online Fraud and AI
We now discuss online fraud and various AI methodologies in detail for further reader information. Online fraud refers to any deliberate act of deception conducted over the Internet to obtain an unlawful or unfair gain. It involves exploiting online platforms, services, and technologies to deceive individuals or organisations for financial, personal, or material gain. Online fraud can take many forms, each characterised by the method of deception and the medium used to perpetrate the fraud.
@@ 1.1 Fraud Categories
The list of online fraud activities is extensive and constantly evolving with new types or sub-types emerging. Developing a comprehensive taxonomy or classification for all online fraud activities requires special attention, which is beyond scope of this work. Below we outline some of the well-known online fraud types.
  • Phishing is the process where fraudsters impersonate representatives of legitimate organisations or acquaintances of the targeted victim to trick them into providing personal information such as usernames, passwords, credit card details, or bank account details. This activity can be done through various mediums, like email, phone calls (aka Vishing), SMS (aka Smishing), and any other way of online communication. There are various phishing scams that surfaced over the years, including the Royal Mail scam royalmailscams , banking scams barcleysscams , HMRC scams hmrcscams , and many others. Notably, phishing scams often include deceptive web addresses created by cyber-criminals to trick victims into believing they are visiting legitimate websites. The primary goal of these URLs is to steal personal data including usernames, passwords and credit card details for financial gain.
  • Fake Reviews are deceptive or fraudulent reviews created to mislead potential customers about the quality, reliability, or legitimacy of a product, service, or app. On fraudulent e-commerce websites and app stores, fake reviews play a crucial role in tricking victims into trusting and using fraudulent apps or purchasing substandard or non-existent products. This leads to potential victims trusting fraudulent websites, services, or apps, providing them with their credit card details for a purchase, that leads the victim to a vulnerable position paul2021fake .
@@
  • Romance fraud (aka romance scams or dating scams) involves fraudsters creating fake profiles on dating websites, social media, or other online platforms to deceive victims into believing they are in a genuine romantic relationship. The primary objective is to exploit the victim’s emotions to extort money, personal information, or other benefits. This is an elaborate scam that is extremely difficult to detect, since it is also underreported due to victims feeling ashamed and hurt for being a victim from someone they considered to be a romantic partner coluccia2020online . In this scams, fraudsters communicate with victims for a long time before presenting them with an “investment opportunity” or requesting their financial aid. Romance scams are closely related to the Cryptocurrency Pig Butchering scams cross2023romance , where victims are gradually lured into making increasing contributions over a long period of time, usually in cryptocurrency, to a fraudulent scheme ordekian2024sinister .
@@
  • Fraudulent Investment include scams where fraudsters promise victims significant winnings or lucrative opportunities vasek2015there . These scams are usually associated with the romance scams discussed above. Once the victims try to withdraw their “winnings” the scammers will extort them by asking for “fees” and “taxes” to be paid in advance. The promised benefits and winnings never materialize, and the initial investment sums and fees are lost agarwal2023defi . Fraudulent investment is the umbrella that covers Cryptocurrency Pig Butchering scams explained above, and various Ponzi schemes vasek2019analyzing where early investors greatly benefit from the investments of later investors, also known as pyramid schemes.
@@
  • Crypto market manipulation involves artificially increasing or decreasing the price of cryptocurrencies to achieve financial gain. It often involves coordinated efforts by individuals or groups to manipulate the market to create false perceptions of supply, demand, or market sentiment. Some common techniques used in crypto market manipulation include: Pump and Dump which inflates the price of a cryptocurrency through misleading or false statements (pumping), encouraging others to buy, and then selling off the cryptocurrency at a profit once the price has been pumped up (dumping); Wash Trading occurs when a trader buys and sells the same cryptocurrency simultaneously to create deceptive activity on the market; Spoofing involves placing significant buy or sell orders to withdrawn them before execution to mislead perceptions related to the market demand or supply;Front-running involves placing orders ahead of a large trade that is known to occur, to benefit from the subsequent price movement caused by the large trade; and many others hamrick2021examination .
@@
  • Fraudulent e-commerce involves deceptive practices or scams conducted through online e-commerce platforms. These scams aim to exploit the digital payment systems to deceive consumers or businesses via paying for a fraudulent product or service.
@@
  • Fraudulent crowdfunding refers to the misuse of crowdfunding platforms to deceive donors or backers, often by providing false or misleading information about the nature, purpose, or outcome of a crowdfunding campaign. Crowdfunding itself is a method of raising money from a large number of people via online platforms to fund projects, products, or causes cumming2021disentangling . A fraud similar to crowdfunding is Charity Fraud and Disaster Scams where scammers seek donations for organisations that do not exist or do no or little work at all. These scams are particularly common after high profile disasters as criminals often use tragedies to exploit people who are looking to donate charityfraud .
@@
  • Gambling Fraud is any illegal activity that is intended to cheat players or an online gambling platform. Fraudsters manage to trick victims and platforms in different ways, including rigged games, fake websites (phishing URLs described above), account takeovers (via stealing legitimate users’ access codes), and via creating fake apps with fake reviews, as discussed above to gain the trust of users. Online gambling fraud can happen on multiple platforms and involve wide variety of games inclduing casino scams, sports betting scams, and lottery scams  hong2022analyzing .
@@
  • Tax Scams occurs when scammers falsify information regarding pending tax money, or via maliciously impersonating tax officials to trick individuals or business entities to willfully pay them ‘fees” brody2014income . Scams similar to tax scams are Council Tax Scams, various Utility bill scams, Insurance Scams, etc., where scammers impersonate officials and trick the victim that they own money. These scams fall under the umbrella of Phishing as they often take place via SMS, phone calls, or emails.
@@
  • Pension Scams are similar to Tax Scams. Scammers aim to make money through fees, direct access to pension savings or by receiving investments pensionscams .
Overall, in this section we briefly described some well known scams and online fraud activities. The complexity and interconnected nature of scams and frauds make it challenging to categorise them under a single typology. Phishing scams, for instance, serve as a broad umbrella that currently encompasses phishing conducting using various methods like vishing (voice) and smishing (sms), yet they can also be integral parts of investment scams when scammers develop phishing websites to gain the trust of victims. Similarly, scams often involve tricking victims through deceptive tactics, ultimately leading to defrauding them of their money or personal information. As such, different scam types frequently overlap, blurring the lines between distinct categories and demonstrating the intricate web of fraudulent activities that exist today. The multifaceted nature of these scams highlights the difficulty in creating a comprehensive classification system that can effectively encompass all types of fraudulent schemes.

### v2

@@ Abstract
Background Fraud is a prevalent offence that extends beyond financial loss, impacting victims emotionally, psychologically, and physically. Advances in online communication technologies continue to create new opportunities for fraud, and fraudsters increasingly using these channels for deception. With the progression of technologies like Generative Artificial Intelligence (GenAI), there is a growing concern that fraud will increase in scale using these advanced methods, with offenders employing deep-fakes in phishing campaigns, for example. However, the application of AI to detect and analyse patterns of online fraud remains understudied. This review addresses this gap by investigating the potential role of AI in analysing online fraud using text data.
@@
Methods We conducted a Systematic Literature Review (SLR) to investigate the application of AI and Natural Language Processing (NLP) techniques for online fraud detection. The review adhered to the PRISMA-ScR protocol, with eligibility criteria including language, publication type, relevance to online fraud, use of text data, and AI methodologies. Out of $2,457$ academic records screened, $350$ met our eligibility criteria, and $223$ were analyzed and included herein.
@@
Results We discuss the state-of-the-art NLP techniques used to analyse various online fraud categories; the data sources used for training the NLP models; the NLP algorithms and models built; and the performance metrics employed for model evaluation. We find that the current state of research on online fraud is broken into the various scam activities that take place, and more specifically, we identify $16$ different frauds that researchers focus on. Finally, we present the most recent and best-performing AI methods employed for detecting online scams and fraud activities.
@@
Conclusions This SLR enhances academic understanding of AI-based detection methods for online fraud and offers insights for policymakers, law enforcement, and businesses on safeguarding against such activities. We conclude that existing approaches focusing on specific scams are unlikely to generalise effectively, as they will require new models to be developed for each fraud type.
Furthermore, we conclude that the evolving nature of scams limits the effectiveness of models trained on outdated data. We also identify that researchers often omit discussions of the limitations of their data or training biases. Finally, we find issues in the consistency with which the performance of models is reported, with some studies selectively presenting metrics, leading to potential biases in model evaluation.
@@
This manuscript has been accepted in the Crime Science Journal. Please cite accordingly.
DOI: 10.1186/s40163-025-00248-8
@@ 1 Introduction
Online fraud has emerged as one of the most pervasive and challenging threats in the digital age, affecting individuals of all ages, businesses of different sizes, and governments. Defined broadly, online fraud is an umbrella term that involves acts of deception or deliberate impersonation on the Internet for the personal gain of the fraudster, often resulting in a financial loss for the victim ukfinance2023 . In addition to financial losses, fraud can have a wide range of impacts on victims. These include emotional and psychological effects such as anger, fear, shame, depression, loss of confidence, and trauma; impacts on physical and mental well-being; it can harm relationships and lead to loneliness and isolation; and cause negative changes in behaviour parliament_post_2020 . Although evidence suggests that certain sociodemographic groups face higher risks of fraud (e.g., women aged 25-44 and those in the highest income bracket), fraud affects individuals across all demographics parliament_post_2020 and sometimes in different ways. For example, in a UK study, victims earning £20,000 or less, those aged 65 and over, and female victims reported that fraud impacted their self-confidence more than did victims in general.
@@
For the year ending March 2023, the Crime Survey for England and Wales estimated that 3.5 million fraud offences, including online fraud, took place that year ons2023 . In that year, compared to the year ending March 2020, advance fee fraud increased significantly, from $60,000$ to $391,000$ offences. This increase is largely due to society’s growing reliance on the Internet and digital platforms for everyday services, transactions, and communications. According to The Office of Communications (Ofcom) ofcom2023 , $92\%$ of adults in the UK use the Internet for a wide variety of activities, including communication, education, and entertainment. Activities such as banking, shopping, and socialising are increasingly happening via online platforms, expanding the landscape for fraudsters to exploit vulnerabilities or use these platforms to deceive victims. In 2020, online shopping scams made up $38\%$ of all reported scams worldwide, an increase of $6\%$ compared to the pre-Covid-19 outbreak statista-ecommerce-fraud .
@@
Online fraud encompasses a wide range of deceptive activities, including identity theft, phishing, advance fee fraud, romance scams, fraudulent investment scams, and more. It is important to highlight that there is no universally accepted definition of “online fraud,” and the term is often used interchangeably with the term “scam”. Legally, “fraud is defined as false representation to cause loss to another or to expose another to a risk of loss” ukfraudlegislation2006 , and scam is the process where criminals gain the trust of victims to deceive or cheat them ukscamglossary2006 through false representation and other means, so that the victim trusts them, which in turn results in various kinds of losses.
The National Fraud Authority of UK published a literature review fraudtypologiesuk , in which they compared the distinction of the term fraud as defined by the amended Fraud Act 2006 ukfraudlegislation2006 and the typology produced by Levi levi2008definingfraud . They found that fraud embraces a broad scope of crimes, whereas scams often focus on fraud against individuals and small firms. For example, different scams like advance fee, romance, tech support, etc., all fall under the fraud umbrella, but they are also deception methods, which are, in part, scams. Hence, in this work, we use both terms as various scams represent the different deception methods scammers use to trick victims, while the term fraud includes all scams.
@@
Online frauds exploit the virtual nature of the Internet and the anonymity it provides to reach victims. This virtual environment, coupled with jurisdictional challenges (where offenders and victims may be in different regions of the world), makes fraud difficult to detect and prevent using traditional policing techniques. The complexity of online fraud is further heightened by its evolving nature, as fraudsters continuously adapt their techniques to bypass new security measures and exploit emerging technologies to target new victims skidmore24 .
@@
Given the scale and impact of online fraud, there is a need for new methods to detect and prevent such activities. The use of Natural Language Processing (NLP), in combination with other Artificial Intelligence (AI), has been proposed for identifying, characterising, and detecting fraudulent patterns in applications like phishing 9795286 , fake job advertisement amaar2022detection , and for the purposes of analysing scam patterns lwin2023supporting which could help develop preventive measures and mitigate risks of online fraud. However, understanding the current state of AI techniques in combating online fraud, the data sources used, the evaluation methods for AI models, and the specific types of fraud that are most prevalent, remains a significant challenge. This is due to the constant emergence of new fraud activities that use various communication mediums and social engineering attacks, in an attempt by fraudsters to remain undetected. Therefore, there is a pressing need to shift from detecting and analysing the effects of fraud to the early detection of emerging fraudulent activities online and new methods of social engineering.
@@
This study aims to address these challenges by conducting a comprehensive review of the state-of-the-art AI techniques used to detect fraudulent online activities. Specifically, we examine the data sources widely used by researchers to study online fraud, the methods researchers use to evaluate the developed AI models, and the most popular types of online fraud targeted in their studies. By synthesising findings from academic papers, this review aims to provide a thorough understanding of the current landscape of online fraud detection and prevention, highlighting gaps in existing research, and proposing directions for future studies.
Manuscript Structure. The rest of the paper is organised as follows. The next section (§2) introduces various well-known types of online fraud and provides a detailed discussion of the latest and most widely used AI methodologies, including how they are evaluated. The background review conducted for this section helped us to formulate and refine our research questions.
Section 3 outlines the methodology followed in this SLR, including the protocol, the criteria used to filter eligible papers, and the data extracted from each study. The results of the SLR are then presented in Section 4. In Section 5, we discuss the findings of our literature review, categorized by the various types of online fraud identified, and provide detailed insights into how each of our research questions were addressed.
Finally, Section 6 offers a deeper analysis of our findings, highlighting limitations and shortcomings in the reporting of AI models, particularly regarding performance and data sources. We also propose recommendations for researchers developing detection models for online fraud, before concluding in Section 7.
@@ 2 Online Fraud and AI
Online fraud refers to any deliberate act of deception conducted over the Internet to cause an unlawful or unfair loss ukfraudlegislation2006 . It involves exploiting online platforms, services, and technologies to deceive individuals or organisations for financial, personal, or material gain. Online fraud can take many forms, each characterised by the method of deception and the medium used.
@@ 2.1 Fraud Categories
The list of online fraud activities is extensive and constantly evolving, with new types and sub-types emerging skidmore24 . To conduct our SLR, it was important first to identify the most prevalent types of offences likely to be analysed by the studies included. To briefly discuss online fraud, we studied various taxonomies, studies, and reports published or discussed by UK government bodies fraudtypologiesuk ; parliament_post_2020 , financial services ukfinance2023 , telecommunication providers ofcom_report , policing think tanks skidmore24 , and academics rabitti2024taxonomy ; zhou2024understanding ; levi2008definingfraud .
Developing a comprehensive taxonomy or classification for all online fraud activities requires special attention, which is beyond the scope of this work. Note that many taxonomies, especially the ones published by UK government bodies fraudtypologiesuk ; parliament_post_2020 also discuss fraud and crime that potentially can take place offline, which we omit in the following discussion of online fraud as offline crime falls beyond the focus of our SLR. Below, we outline some of the well-known and most discussed online fraud types we encountered while performing our preliminary research on online fraud, aided by the reports discussed above. We note that the following is not intended as a complete taxonomy of online fraud, nor does it represent the findings of this SLR. Instead, it is intended to briefly discuss popular online frauds and scams for the reader.
  • Phishing is the process where fraudsters impersonate representatives of legitimate organisations or acquaintances of the targeted victim to trick them into providing personal information such as usernames, passwords, credit card details, or bank account details. This activity can be done through various mediums, like email, phone calls (aka Vishing), SMS (aka Smishing), and any other way of online communication. Various phishing scams have surfaced over the years, including the Royal Mail scam  royalmailscams , banking scams barcleysscams , and HMRC scams hmrcscams . Notably, phishing scams often include deceptive web addresses created by cybercriminals to trick victims into believing they are visiting legitimate websites. The primary goal of these URLs is to steal personal data, including usernames, passwords and credit card details, for financial gain.
  • Fake Reviews are deceptive or fraudulent reviews created to mislead potential customers about the quality, reliability, or legitimacy of a product, service, or app. On fraudulent e-commerce websites and app stores, fake reviews play a crucial role in tricking victims into trusting and using fraudulent apps or purchasing substandard or non-existent products. This leads to potential victims trusting fraudulent websites, services, or apps, providing them with their credit card details for a purchase, which leads the victim to a vulnerable position paul2021fake .
@@
  • Romance Fraud (aka romance scams or dating scams) involves fraudsters creating fake profiles on dating websites, social media, or other online platforms to deceive victims into believing they are in a genuine romantic relationship. The primary objective is to exploit the victim’s emotions to extort money, personal information, or other benefits. This elaborate scam is extremely difficult to detect since it is also under-reported due to victims feeling ashamed and hurt for being victimised by someone they considered to be a romantic partner coluccia2020online . In these scams, fraudsters communicate with victims for a long time before presenting them with an “investment opportunity” or requesting their financial aid. Romance scams are closely related to Cryptocurrency Pig Butchering scams cross2023romance , where victims are gradually lured into making increasing contributions over a long period of time, usually in cryptocurrency, to a fraudulent scheme ordekian2024sinister .
@@
  • Fraudulent Investment includes scams where fraudsters promise victims significant winnings or lucrative opportunities vasek2015there . These scams are usually associated with the romance scams discussed above. Once the victims try to withdraw their “winnings,” the scammers will extort them by asking for “fees” and “taxes” to be paid in advance. The promised benefits and winnings never materialize, and the initial investment sums and fees are lost agarwal2023defi . Fraudulent investment is the umbrella that covers Cryptocurrency Pig Butchering scams explained above, and various Ponzi schemes vasek2019analyzing where early investors greatly benefit from the investments of later investors, also known as pyramid schemes.
@@
  • Crypto Market Manipulation involves artificially increasing or decreasing the price of cryptocurrencies to achieve financial gain. It often involves coordinated efforts by individuals or groups to manipulate the market to create false perceptions of supply, demand, or market sentiment. Some common techniques used in crypto market manipulation include: Pump and Dump, which inflates the price of a cryptocurrency through misleading or false statements (pumping), encouraging others to buy, and then selling off the cryptocurrency at a profit once the price has been pumped up (dumping); Wash Trading occurs when a trader buys and sells the same cryptocurrency simultaneously to create deceptive activity on the market; Spoofing involves placing significant buy or sell orders to withdraw them before execution to mislead perceptions related to the market demand or supply; Front-running involves placing orders ahead of a large trade that is known to occur, to benefit from the subsequent price movement caused by the large trade; and many others hamrick2021examination . These scams are also similar to Stock Market Manipulation.
@@
  • Fraudulent E-Commerce involves deceptive practices or scams conducted through online e-commerce platforms. These scams aim to exploit digital payment systems to deceive consumers or businesses by paying for a fraudulent product or service.
@@
  • Fraudulent Crowdfunding refers to the misuse of crowdfunding platforms to deceive donors or backers, often by providing false or misleading information about a crowdfunding campaign’s nature, purpose, or outcome. Crowdfunding is a method of raising money from many people via online platforms to fund projects, products, or causes cumming2021disentangling . A fraud similar to crowdfunding is Charity Fraud and Disaster Scams, where scammers seek donations for organisations that do not exist or do little work. These scams are particularly common after high-profile disasters as criminals often use tragedies to exploit people who are looking to donate charityfraud .
@@
  • Gambling Fraud is any illegal activity that is intended to cheat players or an online gambling platform. Fraudsters manage to trick victims and platforms in different ways, including rigged games, fake websites (phishing URLs described above), account takeovers (via stealing legitimate users’ access codes), and creating fake apps with fake reviews, as discussed above, to gain the trust of users. Online gambling fraud can happen on multiple platforms and involve a wide variety of games, including casino scams, sports betting scams, and lottery scams hong2022analyzing .
@@
  • Tax Scams occur when scammers falsify information regarding pending tax money or maliciously impersonate tax officials to trick individuals or business entities into wilfully paying them “fees” brody2014income . Scams similar to tax scams are Council Tax Scams, various Utility bill scams, Insurance Scams, etc. These scams fall under the umbrella of Phishing as they often take place via SMS, phone calls, or emails.
@@
  • Pension Scams are similar to Tax Scams. Scammers aim to make money through fees, direct access to pension savings, or by receiving investments pensionscams .
The complexity and interconnected nature of scams and frauds make categorising them under a single typology challenging skidmore24 ; cohen2019investigation . Phishing scams, for instance, serve as a broad umbrella that covers phishing conducted using various methods like vishing (voice) and smishing (SMS). Yet, they can also be integral parts of investment scams when scammers develop phishing websites to gain victims’ trust. Similarly, most scams often involve the scammer impersonating an authority (government, law enforcement), friend, organisation (e.g., bank), or other entity (e.g., delivery service), making impersonation scams difficult to break down as they are an integral part of other scams (e.g., a delivery scam is also an impersonation scam as the offender is clearly not an Amazon representative, for example).
To summarise, different scam types frequently overlap, blurring the lines between distinct categories and demonstrating today’s intricate web of fraudulent activities. The multifaceted nature of these scams highlights the difficulty in creating a comprehensive classification system that can effectively encompass all types of fraudulent schemes. We discuss this challenge and limitation in detail before concluding our SLR, in our Discussion (Section 6).